OpenStudy (anonymous):

Is 2x+3y=6 parallel to 2x-3y=6

OpenStudy (anonymous):

solve them both for "y" and plug in -1, 0, and 1 into both and see.

OpenStudy (anonymous):

When we can say two line are parallel they have equal slopes(m) in standard form . Standard form is the form we say for the line and this a a equation between Y and X of the line . It s : Y=mx+d m is slope . It s before x . y dont have any number in its after . So if we have should divided the equation on that number . And x and y are not together . We have both of this there so divided and change the places . :)

OpenStudy (anonymous):

we should get 2x+3y=6 in standard form . So first have y in a place and x in another place . Means : 2x+3y=6 => 3y = -2x +6 Then divided on 3 : y=-2/3x+2 So we have a equation in standard form now . And for : 2x-3y=6 do this works to .

OpenStudy (anonymous):

we have : 2x-3y=6 => -3y=-2x+6 divided on -3 : y=2/3x-2 What is the slopes ? @gunterlove

OpenStudy (anonymous):

in my class we are not so complex in doing these type of problems im in 8th grade alegbra

OpenStudy (anonymous):

We have tow equation now : y=2/3x-2 => m = 2/3 y=-2/3x+2 => m = -2/3 Is -2/3 = 2/3 ? No ! so we cant say they are parallel .;) Got it ?:) @gunterlove By the way ... Welcome to open study !;)

OpenStudy (anonymous):

so the answer would be 2x+3y=6 is parallel to 2x+3y=-6

OpenStudy (anonymous):

No ! You forgot standard form ! See my answer again ;)

OpenStudy (anonymous):

Oh i get it now thank you so much for all your help have a good rest of the day!

OpenStudy (anonymous):

;) You re welcome ! I wish my answer could help you ;)

OpenStudy (anonymous):

which property would -5x+3+2x>9 be

OpenStudy (anonymous):

what is the x-intercept of 3x-4y-12=0

OpenStudy (anonymous):

factor, set factor =0 to solve for x intercepts.